Shift summary for day shift 8/19/99

Taku Yamanaka (taku@fnal.gov)
Thu, 19 Aug 1999 15:44:46 -0500

Shift Summary for eve shift, 1999, August 16.

On Shift: ER, TY

Hours of beam available at all intensities:
Hours of beam available at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

========================================================================
Runs ended this shift:
Run 14272 0.8M events - BEAM_832 beam has been wondering around
Run 14273 3.4M events - BEAM_832
Run 14275 1.7M events - BEAM_832

Tapes brought to FCC this shift:
(including tapes for online split)

QKC285-QKC287
QKS285-QKS287
QKK285-QKK287
QKB100

========================================================================
Accesses during shift:

- None

========================================================================
Overview of shift:
E832 datataking

========================================================================
Description of Shift:

Beam position has been moving wildly till ~8:25, but fixed by RickF.
Otherwise, more or less a smooth data taking.

Intensity is around 5.5 - 6E12, a little lower than request, due to
lower intensity in the ring.

CsI ch 226: gain is low by 5%.

AC is fixed.

Taku's first shift this year. Rehabilitated to do kfrend, run
kumacs,...

run spill time comment
------------------------------------------------------------------------
14272
7:00 Shift starts.
Beam position is moving around.
7:35 RickF noticed that NM2V is not ramping.
8:12 Stopped Run.

14273 8:27 Started Run. BEAM_832. Beam is now stable.
9:45 Laser crate BUSY does not go away. Power cycled
the crate
Crate 3-53 BUSY does not go away, Power cycled it,
too.
10:00 Found that spill numbers stayed at zero.
Stopped Run.

14274 10:10 Starting Run failed. fresh_start.

Online Split asked to remount two tapes in
ktev4 and ktev2.

14275 10:30 Started Run. BEAM_832
10:31 Beam down for testing NRF safety interlock.
11:13 Beam is back.

11:50 Linac RF station problem, EST 20min.
12:10 Stopped Run, since we lost info on tape usage.

14276 12:25 Start Run (BEAM_832) on new tapes,
QK*288-290,QKB101.
15:02 Beam is turned off.
15:27 Beam is back, but NM2 H position is off. Ops is
tuning it.
15:38 LB is here, but not Brad.
15:39 Brad is here
15:40 AC is fixed. They replaced a loose fan belt, and
a faulty temperature sensor.

========================================================================
Follow up Notes on Previous Shift

CsI ch.226: gain is still low by 5%.
beam position: Fixed by paging RickFord. He is the man.

========================================================================
Items to Follow up on Next Shift

CsI ch.226: If it gets larger than 5% off, call Sydney immediately.
Sydney: If you get a problem, call him. He will be happier to be
called than not being called as last night.